OpenStudy (anonymous):

Express as a single logarithm, and, if possible, simply. 1/4 loga x + 6 loga y-2loga x

OpenStudy (anonymous):

Hint: xlogy =log(y^x) logx+logy = log(xy)

OpenStudy (anonymous):

Does it help?

OpenStudy (anonymous):

No because I am clueless when it comes to logs

OpenStudy (anonymous):